soybean is 12% protein, cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240 pounds of mixture that is 7% protein?
Answer by stanbon(75887)

soybean is 12% protein, cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240 pounds of mixture that is 7% protein?
--------
Equation:
protein + protein = protein
0.12x + 0.06(240-x) = 0.07*240
12x + 6*240 - 6x = 7*240
6x = 240
x = 40 lbs (amt. of soybeen needed)
---
240-x = 200 lbs (amt. cornmeal needed)
=======================================
Cheers,
Stan H.
=============